OpenStudy (anonymous):

can someone please help explain.. how to simplify and write in exact radical form √15x^3 (times) √3x^2

OpenStudy (anonymous):

@jdoe0001 do you get how to do this?

OpenStudy (jdoe0001):

one sec

OpenStudy (anonymous):

okay

OpenStudy (jdoe0001):

\(\bf \sqrt{15x^3}\cdot \sqrt{3x^2}\implies \sqrt{15\cdot 3\cdot x^{\color{brown}{ 3}}\cdot x^{\color{brown}{ 2}}}\implies \sqrt{45\cdot x^{{\color{brown}{ 3+2}}}} \\ \quad \\ {\color{blue}{ 45\to 3\cdot 3\cdot 5\to 3^2\cdot 5}} \qquad thus \\ \quad \\ \sqrt{45\cdot x^{{\color{brown}{ 3+2}}}}\implies \sqrt{3^2\cdot 5\cdot x^5}\implies 3\sqrt{5x^5} \\ \quad \\ {\color{blue}{ x^5\to x^{2+2+1}\to x^2\cdot x^2\cdot x^1}}\qquad thus \\ \quad \\ 3\sqrt{5x^5}\implies 3\sqrt{5x^2x^2x}\implies 3\cdot x\cdot x\sqrt{5x}\implies 3x^2\sqrt{5x}\)

OpenStudy (anonymous):

alright I see how you got the answer but I am unsure about your 3rd line of work where you break up the exponent. Could you explain it ?

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

well... tis a simple exponent rule \(\bf a^n\cdot a^m\implies a^{n+m}\) and from there you can use the "commutative property" so... in say... for example if we had.... \(\large a^3\cdot a^{11}\implies a^{3+11}\implies a^{14} \begin{cases} a^{9+5}\to a^9\cdot a^5\\ a^{10+4}\to a^{10}\cdot a^4\\ a^{7+7}\to a^7\cdot a^7\\ a^{20-6}\to a^{10}\cdot a^{-6}\\ \textit{and so on ....} \end{cases}\)

OpenStudy (jdoe0001):

hmmm well \(\large a^3\cdot a^{11}\implies a^{3+11}\implies a^{14} \begin{cases} a^{9+5}\to a^9\cdot a^5\\ a^{10+4}\to a^{10}\cdot a^4\\ a^{7+7}\to a^7\cdot a^7\\ a^{20-6}\to a^{20}\cdot a^{-6}\\ \textit{and so on ....} \end{cases}\)

OpenStudy (jdoe0001):

in the case above, you had \(x^5\) and the root is square, or root 2 so we took out all "2" possibles out of 5 so 2 + 2 + 1 and those come out of the radical we pretty much did the same with 45 if you notice broke down to 3 * 3 * 5 so we took out all possible squares, thus \(3^2\)

OpenStudy (anonymous):

thank you so much!

OpenStudy (jdoe0001):

yw